OpenStudy (anonymous):

in a recent year the total sales at resterraunt A were $100,000 more than resteraunt B total sales at the two stores were $575,000,000. what were the sales for the year at each store?

OpenStudy (anonymous):

A = B + 100,000 A + B = 575,000 now we sub B + 100,000 in for A in the 2nd equation A + B = 575,000 (B + 100,000) + B = 575,000 -- combine like terms 2B + 100,000 = 575,000 -- subtract 100,000 from both sides 2B = 575,000 - 100,000 2B = 475,000 -- divide both sides by 2 B = 475,000/2 B = 237,500 A = B + 100,000 A = 237,500 + 100,000 A = 337,500 check.. A + B = 575,000 337,500 + 237,500 = 575,000 575,000 = 575,000 (correct) Store A sales for the year are $337,500 Sore B sales for the year are $ 237,500